In a study of 50 new cars, 17 of them were white. Find p ̂ and q ̂ where p-hat is the proportion of new cars that is white.
Answer by stanbon(75887)

In a study of 50 new cars, 17 of them were white. Find p ̂ and q ̂ where p-hat is the proportion of new cars that is white.
----
p-hat = 17/50 = 0.34
----
q-hat = 33/50 = 0.66
--------------------------
cheers,
Stan H.
=========================